A vertebra is subjected to a shearing force of 686 N. Find the shear deformation, taking the vertebra to be a cylinder 3.29 cm high and 7.69 cm in diameter


1
Expert's answer
2021-03-29T09:07:03-0400

The shear deformation is


"\delta_s=\frac{FH}{AG}=\frac{4FH}{\pi d^2G}=\\\space\\
\delta_s=\frac{4·686·0.0329}{\pi·0.0769^2·8·10^{10}}=6.07·10^{-8}\text{ m}."